# Promotion of Scientific Publications on ArXiv and X Is on the Rise and Impacts Citations

### v1

@@ Abstract
In the evolving landscape of scientific publishing, it is important to understand the drivers of high-impact research, to equip scientists with actionable strategies to enhance the reach of their work, and to understand trends in the use modern scientific publishing tools to inform their further development. Here, based on a large dataset of computer science publications, we study trends in the use of early preprint publications and revisions on ArXiv and the use of X (formerly Twitter) for promotion of such papers in the last 10 years. We find that early submission to ArXiv and promotion on X have soared in recent years. Estimating the effect that the use of each of these modern affordances has on the number of citations of scientific publications, we find that in the first 5 years from an initial publication peer-reviewed conference papers submitted early to ArXiv gain on average $21.1\pm 17.4$ more citations, revised on ArXiv gain $18.4\pm 17.6$ more citations, and promoted on X gain $44.4\pm 8$ more citations. Our results show that promoting one’s work on ArXiv or X has a large impact on the number of citations, as well as the number of influential citations computed by Semantic Scholar, and thereby on the career of researchers. We discuss the far-reaching implications of these findings for future scientific publishing systems and measures of scientific impact.
@@ Introduction
Scientific publications started in the $17^{th}$ century when decisions about whether a paper would be published were made by a select group of experts. As scientific publications proliferated in the $20^{th}$ century, editors of these journals alone could no longer make informed decisions about publications for such a large volume of papers. This led to peer review becoming a standard practice in academia. In the 1990s, the advent of the internet greatly helped with the dissemination of scientific publications. Modern science offers affordances such as publishing preprints before peer-review, revising preprints, promoting scientific works on social media, and viewing researcher profiles listing all co-authored publications and the overall number of citations. However, it is not clear what are the trends in scientific publishing and evaluation, what is their impact, and what will the future bring.
Open-access e-print repositories such as ArXiv and SSRN, and publishing platforms such as OpenReview.net and PeerJ are at the forefront of this change (Sutton and Gong 2017; Abdill and Blekhman 2019), facilitating swift exchange of scientific information. ArXiv hosts 2 million e-prints in various technical fields, while SSRN houses 1.3 million papers for social sciences. These platforms enable early publications and later revisions of unreviewed manuscripts. Then, once a preprint is published, the work can be instantly promoted on social media. Early preprint publication and promotion on social media likely contribute to the number of citations of respective papers, which arguably is the most popular measure of scientific impact. This phenomenon incentivizes scientists to publish early preprints and promote them on social media, without any peer review of their works, suggesting that the publication of preprints and promotion on social media will become standard practices, if not the case already.
Unfortunately, these modern affordances of scientific publishing process, can lead to the spread of questionable scientific results. For instance, in July 2023, a Korean team claimed in two ArXiv publications the discovery of a room-temperature superconductor, which quickly became widely discussed on social media. By mid-August a series of papers from major labs debunked the claim (Johnson 2023). Prior works suggest that social influence make it more challenging to distinguish between high and low quality content (Salganik, Dodds, and Watts 2006), including in the context of scientific publications (Fisher and Parisis 2015; Baddeley 2015; Resnik and Smith 2020), and recent research introducing novel measures of scientific impact discusses a correction for such social impact (Ke, Gates, and Barabási 2023).
@@
Figure 1: Top-30 computer science conferences by paper count in our cross-referenced dataset. The colors correspond to research areas in computer science according to Sutton and Gong (2017).
Despite these potential implications of early preprint publications and their promotion on social media, there is a lack of comprehensive, large-scale studies on the use of these affordances and their impact on the number of citations. This study bridges this gap and reflects on the future of scientific publishing. The study focuses on the field of computer science due to its rapid adoption of new scientific publishing technologies. In the field of computer science, ArXiv and X are currently the go-to platforms for publishing preprints and disseminating research findings. Specifically, we address the following research questions.
@@
ArXiv offers next-day publication that can lead to a first-mover advantage and boasts of over 5 million monthly active users (Tech 2023). However, it is unclear whether publishing first on ArXiv, rather in a conference proceeding, impacts the number of citations.
@@
Finally, we hypothesize that more modern measures of scientific impact may be less susceptible to influence of social media, following the intuition that important citations correspond to a deep semantic connection between the citing and cited papers.
@@
These research questions have not been sufficiently addressed yet, since it is challenging to answer them quantitatively. Performing a controlled experiment across multiple conferences over multiple years through intervention is intractable. Calculating treatment effects using observational data is difficult because it requires bridging mutliple datasets. To this end, we need information about peer-reviewed conference publications, ArXiv publications, relevant posts on X, and fine-grained information about the dates of publications. Another challenge is that there are multiple factors that could confound treatment effect estimates.
In this paper, we study publications in computer science conference proceedings across multiple areas. Computer science has a unique publication system, prioritizing conferences over journals. Precise publication dates are crucial for measuring trends in early preprint submissions to ArXiv (RQ1) and for estimating its effect (RQ2). While we utilized WikiCFP to gather conference dates, the same level of precision was not attainable for journal publications, which is the prominent mode of publication outside of computer science. Historically, the field of computer science has been relatively quick to embrace emerging technologies for scientific publication, facilitating the creation of datasets necessary for conducting this study, such as WikiCFP. To answer the research questions addressed in this paper, we create a dataset by cross-referencing five data sources: Semantic Scholar, ArXiv, WikiCFP, Altmetric.com, and X, which includes the largest and most prestigious computer science venues (Fig. 1). The dataset contains 18113 papers published over 34 years.
@@
To account for confounding biases in the dataset, we rely on treatment effect estimation techniques, including state-of-the-art doubly robust estimators. We define treatments that address our four main research questions (RQ2-5) and study their effects on the number of citations of a paper (RQ2-4), or influential citations (RQ5), in the 5 years since the date of first publication (on ArXiv or conference proceedings). We use the terms ArXiv-first effect (RQ2), revision effect (RQ3), effect of tweeting (RQ4.1), and effect of author(s) tweeting (RQ4.2) to refer to these effects in the remainder of this paper. We estimate the effects while controlling for potential confounders, such as longer time to accrue citations and self-selection of better papers in hot research areas to be posted on ArXiv before conference publication.
Our study has implications for future publishing models and measures of scientific impact, as we discuss in the last section of this manuscript, which are particularly relevant given that the computer science community is uniquely positioned to lead the development of novel publishing models and scientific impact measures.
@@
While extensive literature has studied characteristics of open access publications, described next, in this study we investigate a distinct set of questions. Our inquiry centers on the impact of submitting earlier to ArXiv than to a peer-reviewed venue, and subsequently revising. Additionally, we investigate the effect of promoting research on social media. While Feldman, Lo, and Ammar (2018) have explored the ArXiv-first effect (RQ2) in a smaller study without doubly robust estimation, the effects outlined in RQ1 and RQ3-5 remain unexplored or underexplored.
@@
Publishing on ArXiv and promoting a paper on X have increased in popularity in the computer science community. Sutton and Gong (2017) observed a remarkable surge in ArXiv publishing within computer science, reporting that 23% of all papers in 2017 were published on ArXiv (among which 56% were preprints), compared to just 1% in 2007. We study trends in early ArXiv submissions, revisions, and social media promotion for computer science publications from 2013 to 2022.
@@
We measure the effects of submitting a version of one’s paper to ArXiv before the conference, i.e., the ArXiv-first effect. Elazar et al. (2023) also studied early ArXiving but focused on its effect on conference acceptance rather than citations which is our focus. Feldman, Lo, and Ammar (2018), on the other hand, performed an analysis on citations and found that there is a 65% increase in citations that can be attributed to the ArXiv-first effect. This study, however, analyzed only the top 16 conferences, used a subset of the confounders we take into account, and did not apply doubly robust estimation.
Multiple prior works find that open access (OA) papers have higher number of citations (Harnad and Brody 2004; Antelman 2004; Cheng and Ren 2008; Fu and Hughey 2019; Alkhawtani, Kwee, and Kwee 2020). This relation was thought to be causal until a study by Kurtz et al. (2005) about research articles in astrophysics on ArXiv showed that this difference in the number of citations can be explained by two sources of bias that were previously unaccounted for. First, the early view bias refers to the situation where articles that are submitted early to a preprint repository, before peer-review publications, have more time to accrue citations. Second, the self-selection bias refers to the phenomenon that better papers are selected to be published in preprint repositories in additoin to peer-reviewed venues. Moed (2007) states that the self-selection bias can occur due to two reasons: i) the author is very well-reputed and is confident about putting out unfinished work (established author bias), ii) the author, irrespective of their reputation in the field, publish their better works on ArXiv (quality bias). These papers in some cases can be unfinished and the authors keep revising them with time. Controlling for these two biases makes the OA advantage negligible for papers in astrophysics. Since then, there have been multiple studies in different disciplines that study these effects. Davis and Fromerth (2007) find that for mathematics papers there is no significant early view effect and the difference in citations between OA and non-OA can be explained by the self-selection bias. More specifically, they find that there is a quality differential between papers that get submitted to ArXiv compared to papers that do not.
@@
We explore the impact of paper revisions on citations (termed the ”revision effect”). The prior research on this topic is scarce. Rigby, Cox, and Julian (2018) examined the correlation between revisions and the number of citations for a single social sciences journal. We perform a study at a much larger scale for computer science publications and use treatment estimation techniques to account for confounders.
@@
The effect of tweeting remains unexplored in computer science. We quantify the impact of tweeting on the number of citations and examine the impact of authors tweeting about their work versus tweets from other accounts.
@@
We create our dataset using data from five sources (as depicted in Fig. 2). In this section, we describe the details of our dataset development process. Then, we describe the distribution of the number of citations of papers belonging to that dataset and how we use the dataset to answer our research questions.
@@
We have created our dataset using data from five sources: Semantic Scholar 111www.semanticscholar.org (CC BY-NC license), WikiCFP 222www.wikicfp.com, ArXiv 333arxiv.org (CC0: Public Domain license), Altmetric.com, and the academic X API. Semantic Scholar is a search engine for scientific literature developed by the Allen Institute of AI. We used their publicly available dataset to get information about articles, authors, citations, and highly influential citations. WikiCFP is a website that contains details about calls for papers for scientific and technological conferences. We use the data from WikiCFP as a source of important dates relating to the submission and decision deadlines for a conference. To access information about paper submission and revision dates, we utilize the dataset provided by ArXiv. We examined citations spanning a five-year timeframe from the date of first publication (either on ArXiv or in a conference), incorporating papers up to 2017 in our analysis  [^4] Association for Computational Linguistic’s (ACL) anonymity policy started in October 2017 so we have safely ignored that and have not taken the anonymity policy into account in our analysis., as we used Semantic Scholar data up to 2022. Altmetric.com provided us with research access to their proprietary dataset that serves as the data and metadata source for tweets related to each research paper in our dataset. Utilizing the unique tweet ID provided by Altmetric.com, we retrieve the tweet, its metadata, and the author’s metadata using the X API.
@@
Cross-referencing WikiCFP using conference names.
We use WikiCFP as a source of the following 5 dates: the submission deadline, the notification date (the date when the author is notified about acceptance or rejection), final version (camera-ready) due date, and conference start and end dates. Some important conferences in computer science (like NeurIPS, AAAI, and ACL conferences) did not have all the dates on WikiCFP. We gathered all conferences from CSRankings.org and found that 8 of the top conferences  [^5] The names of these conferences can be found in the Appendix B section. were missing from WikiCFP which we added to our dataset. In addition to these 8 conferences, we identified around 400 conference editions with missing or incomplete WikiCFP information. We annotated their dates manually, using information publicly available on the internet, to include as many important computer science venues as possible and to increase our dataset size. Despite this, we could not find all 5 required dates for around 150 important conferences, e.g., some of the older NeurIPS editions. Overall, we have 46.5k conference editions with full information about them. This data is included in our dataset release  [^6] Code and dataset release: http://bit.ly/48OSnE1.
We matched conference names between the WikiCFP and Semantic Scholar datasets to add information about conference dates for the articles on Semantic Scholar. We noticed a lot of variability in conference names between the two datasets as there are no naming conventions that are strictly followed in either of the two datasets. We extended the conference name matching algorithm used by Demetrescu et al. (2022), but unlike their approach, we do not rely on DBLP for conference names. Instead, we utilize conference names and acronyms from WikiCFP , and conference names from Semantic Scholar. Appendix C contains the details of our algorithm. Our final dataset after cross-referencing Semantic Scholar, WikiCFP, and ArXiv contains 18113 articles with author and citation information (including highly influential citations), conference dates, and ArXiv submission data and is used for calculating trends for ArXiv-first submissions and revisions (RQ1) and to calculate the ArXiv-first and revision effects (RQ2,3).
Figure 3: Comparison of the probability distribution for citations in 5 years after the first date of publication for all computer science papers on Semantic Scholar (SS), those on ArXiv cross-referenced with Semantic Scholar, and our dataset derived from cross-referencing ArXiv, Semantic Scholar, and WikiCFP.
@@
The Altmetric.com API provides rich information about papers that are publicized on social media, including aggregate information (for e.g., total number of posts about a paper on social media), as well as more fine-grained information about each social media post describing a paper. Using the API, we found 13564 papers that have at least one tweet associated with them, 2929 papers that have no tweets associated with them, and Altmetric.com did not contain information about 1620 papers in our dataset.
@@
  • Identifying aggregator accounts. To identify aggregators, we set a threshold based on the frequency of tweeting research papers after a close manual inspection, classifying accounts exceeding one paper every 10 days as aggregators. We integrate data on these aggregators’ tweets and followers as a confounding variable in our analysis.
  • Matching authors names to X profiles. We identify papers that authors tweeted about by cross-referencing author names with X users who tweeted about a paper. We employ a name matching algorithm and estimate false negatives using an independent validation classifier. Our name matching algorithm identifies 1,344 papers tweeted by authors, with a true positive rate of 94%. Employing a method from Mane, Srivastava, and Hwang (2005), we estimated approximately 170 false negatives, constituting about 1% of papers in our dataset, which we conclude is negligible. Appendix D contains details on the algorithm and false negative estimation.
@@
The dataset we released contains 18113 articles with author and citation information (including highly influential citations), conference dates, and ArXiv submission data. This cross-referenced dataset has a different citation distribution with a higher number of mean and median citations compared to all computer science papers in Semantic Scholar or ArXiv, since only the higher-tier conferences appear in WikiCFP (compare red and green lines in Fig. 3). Our estimates of treatment effects are for papers that are published in such high-tier conferences. We note that for lower-tier conferences, treatment effects will be proportionally lower. Finally, we notice that the number of influential citations is about 7 times smaller than the number of plain citations (compare dashed green and continuous lines in Fig. 3).
Dataset use for treatment estimation
We use all 18113 papers to study the ArXiv-first effect (RQ2) and revision effect (RQ3), and 16493 (=13564+2929) papers for which we can extract tweet information from Altmetric.com to study the effect of tweeting (RQ4.1). To examine the impact of author(s) tweeting as opposed to receiving tweets from a different source (RQ4.2), we focus on the subset of 13564 papers that have associated tweets. In RQ5, we use the same dataset for confounders and treatment variables as in RQ2-4, with highly influential citations serving as the outcome variable.
@@
This section describes the control variables utilized in our study and the rationale behind using them, followed by details about feature selection and the models used for treatment effect estimation.
Estimates of effects rely on model specification. If model specification is correct, then estimates tend to be unbiased. Doubly robust methods estimate both the model of treatment and outcome. If one of the models is correct, that is, it matches reality, then the estimate is unbiased. Here, to increase the probability that our estimates of average treatment effects are unbiased, we use doubly-robust methods such as Targeted Maximum Likelihood Estimation (TMLE) (Schuler and Rose 2017). Furthermore, we use multiple model specifications and train a superlearner (Van der Laan, Polley, and Hubbard 2007) for our treatment and outcome models. In addition, to capture the distribution of treatment effects among the papers, we perform standardization using the superlearner outcome model.
Confounding Variable Reason to Control
@@
Conference rating

### v2

# Effects of Research Paper Promotion via ArXiv and X

@@ Abstract
In the evolving landscape of scientific publishing, it is important to understand the drivers of high-impact research, to equip scientists with actionable strategies to enhance the reach of their work, and to understand trends in the use of modern scientific publishing tools to inform their further development. Here, based on a dataset of over 0.5 million publications in computer science and physics, we study trends in the use of early preprint publications and revisions on ArXiv and the use of X (formerly Twitter) for promotion of such papers. We find that early submissions to ArXiv and promotion on X have soared in recent years. Estimating the effect that the use of each of these modern affordances has on the number of citations of scientific publications, we find that peer-reviewed conference papers in computer science that are submitted early to ArXiv gain on average $21.1\pm 17.4$ more citations, revised on ArXiv gain $18.4\pm 17.6$ more citations, and promoted on X gain $44.4\pm 8$ more citations in the first 5 years from an initial publication. In contrast, journal articles in physics experience comparatively lower boosts in citation counts, with increases of $3.9\pm 1.1$, $4.3\pm 0.9$, and $6.9\pm 3.5$ citations respectively for the same interventions. Our results show that promoting one’s work on ArXiv or X has a large impact on the number of citations, as well as the number of influential citations computed by Semantic Scholar, and thereby on the career of researchers. These effects are present also for publications in physics, but they are relatively smaller. The larger relative effect sizes, effects of promotion acumulating over time, and elevated unpredictability of the number of citations in computer science than in physics suggest a greater role of world-of-mouth spreading in computer science than in physics. We discuss the far-reaching implications of these findings for future scientific publishing systems and measures of scientific impact.
@@ Introduction
Scientific publications started in the $17^{th}$ century when decisions about whether a paper would be published were made by a select group of experts. As scientific publications proliferated in the $20^{th}$ century, editors of these journals alone could no longer make informed decisions about publications for such a large volume of papers. This led to peer review becoming a standard practice in academia. In the 1990s, the advent of the Web greatly helped with the dissemination of scientific publications. Modern science offers affordances such as publishing preprints before peer-review, revising preprints, promoting scientific works on social media, and viewing researcher profiles listing all co-authored publications and the overall number of citations. However, it is not clear whether the use of these affordances is on the rise and what is their impact.
Open-access e-print repositories such as ArXiv, which was founded by a physicist, and open publishing platforms such as OpenReview.net, founded by a computer scientist, are at the forefront of this change (Sutton and Gong 2017), facilitating swift exchange of scientific information. ArXiv hosts 2 million e-prints in various technical fields, including computer science and physics. These platforms enable early publications and later revisions of unreviewed manuscripts. Then, once a preprint is published, the work can be instantly promoted on social media. Early preprint publication and promotion on social media likely contribute to the number of citations of respective papers, which arguably is the most popular measure of scientific impact. This phenomenon incentivizes scientists to publish early preprints and promote them on social media, without proper peer review of their works, suggesting that the publication of preprints and promotion on social media will become standard practices, if not the case already.
Unfortunately, these modern affordances of scientific publishing process, can lead to the spread of questionable scientific results. An instance of this involves COVID-19 studies published on preprint repositories prior to peer-review that were later retracted due to ethical concerns or research misconduct. These publications still managed to garner up to 593 citations, thereby disseminating misinformation (Syed et al. 2023). Many of these studies circulated on platforms such as Reddit and 4chan, where their findings were amplified, and premature results were frequently misinterpreted to make future predictions (Yudhoatmojo, De Cristofaro, and Blackburn 2023). Additionally, prior works suggest that social influence make it more challenging to distinguish between high and low quality content (Salganik, Dodds, and Watts 2006), including in the context of scientific publications (Fisher and Parisis 2015; Baddeley 2015; Resnik and Smith 2020), and recent research introducing novel measures of scientific impact discusses a correction for such social impact (Ke, Gates, and Barabási 2023).
@@
Figure 1: (a) Top-30 computer science conferences by paper count in our cross-referenced dataset. The colors correspond to research areas in computer science according to Sutton and Gong (2017). (b) Top-10 physics journals by paper count. “PR” refers to Physical Review and “CQG” referes to Classical and Quantum Gravity.
Despite these potential implications of early preprint publications and their promotion on social media, there is a lack of comprehensive, large-scale studies on the use of these affordances and their impact on the number of citations. This study bridges this gap and reflects on the future of scientific publishing. The study focuses on the fields of computer science and physics due to their rapid adoption of new scientific publishing technologies. We address the following research questions.
@@
ArXiv offers next-day publication that can lead to a first-mover advantage and boasts of over 5 million monthly active users (ArXiv 2023). Publishing first on ArXiv, rather than in a peer-reviewed venue, may impact the number of citations.
@@
Next, we hypothesize that more modern measures of scientific impact may be less susceptible to influence of social media, following the intuition that important citations correspond to a deep semantic connection between the citing and cited papers.
@@
Finally, we study the differences in treatment effects between publications in computer science and physics, while aiming to understand the underlying mechanisms that might contribute to these differences.
RQ6: How are the effects of social promotion different for computer science compared to physics?
@@
These research questions have not been sufficiently addressed yet, since it is challenging to answer them quantitatively. Performing a controlled experiment across multiple venues over multiple years through intervention is intractable. Calculating treatment effects using observational data is difficult because it requires bridging mutliple datasets. To this end, we need information about peer-reviewed publications, ArXiv publications, relevant posts on X, and fine-grained information about the dates of publications and citations. Another challenge is that there are multiple factors that could confound treatment effect estimates.
In this paper, we investigate publications in computer science and physics. While computer science publications predominantly occur through conference proceedings, physics publications primarily take the form of journal articles. Both fields contribute significantly to ArXiv. However, the submission rate of computer science papers to ArXiv has drastically increased at a rate 3.5 times higher than that of physics over the last decade (ArXiv 2024).
To answer these research questions, we cross-reference six data sources: Semantic Scholar, ArXiv, WikiCFP, Crossref, Altmetric.com, and X. The resulting dataset includes the largest and most prestigious computer science (Fig. 1(a)) and physics (Fig. 1(b)) venues, such as ICML and PRL, respectively. The dataset includes 18,113 computer science and 501,766 physics papers spanning 34 years. We release this dataset publicly, including author and citation information, publication dates, and ArXiv submission data. [^1] Code and dataset release: http://bit.ly/48OSnE1
@@
We define treatments that address our four main research questions (RQ2-5) and study their effects on the number of citations of a paper (RQ2-4), or influential citations (RQ5), in the 5 years since the date of first publication (on ArXiv or conference proceedings). We use the terms ArXiv-first effect (RQ2), revision effect (RQ3), effect of tweeting (RQ4.1), and effect of author(s) tweeting (RQ4.2) to refer to these effects in the remainder of this paper. To estimate treatment effects, we rely on state-of-the-art doubly robust estimation. We estimate the effects while controlling for potential confounders, such as longer time to accrue citations and self-selection of better papers in hot research areas to be posted on ArXiv before publishing in a peer-reviewed venue. Finally, we compare the measured effects across publications in computer science and physics (RQ6).
We recognize the limitations of using citations as a sole metric for assessing scientific impact. For example, papers with erroneous findings can still accumulate numerous citations as researchers cite them to correct the record. Our study demonstrates that citations are not only susceptible to such inherent flaws, but also highly influenced by social promotion. Our results have implications for future publishing models and metrics for gauging scientific impact, as we discuss in the last section of this manuscript, which are particularly relevant given that the computer science and physics communities are uniquely positioned to lead the development of novel publishing models (as exemplified by ArXiv and OpenReview.net) and scientific impact measures (Ke, Gates, and Barabási 2023).
@@
While extensive literature has studied characteristics of open access publications, described next, in this study we investigate a distinct set of questions. Our inquiry centers on the impact of submitting earlier to ArXiv than to a peer-reviewed venue, and subsequently revising. Additionally, we investigate the effect of promoting research on social media. While Feldman, Lo, and Ammar (2018) have explored the ArXiv-first effect (RQ2) in a smaller study without doubly robust estimation, the effects outlined in RQ1 and RQ3-6 remain unexplored or underexplored.
@@
Publishing on ArXiv and promoting a paper on X have increased in popularity in the computer science community. Sutton and Gong (2017) observed a remarkable surge in ArXiv publishing within computer science, reporting that 23% of all papers in 2017 were published on ArXiv (among which 56% were preprints), compared to just 1% in 2007. We study trends in early ArXiv submissions, revisions, and social media promotion for computer science and physics publications from 2013 to 2022.
@@
We measure the effects of submitting a version of one’s paper to ArXiv before the conference, i.e., the ArXiv-first effect. Elazar et al. (2023) also studied early ArXiving but focused on its effect on conference acceptance rather than citations which is our focus. Feldman, Lo, and Ammar (2018), on the other hand, performed an analysis on citations and found that there is a 65% increase in citations that can be attributed to the ArXiv-first effect. This study, however, analyzed only the top 16 conferences in computer science, used a subset of the control variables we take into account, and did not apply doubly robust estimation.
Multiple prior works find that open access (OA) papers have higher number of citations (Harnad and Brody 2004; Antelman 2004; Fu and Hughey 2019; Alkhawtani, Kwee, and Kwee 2020). This relation was thought to be causal until a study by Kurtz et al. (2005) about research articles in astrophysics on ArXiv showed that this difference in the number of citations can be explained by two sources of bias that were previously unaccounted for. First, the early view bias refers to the situation where articles that are submitted early to a preprint repository, before peer-review publications, have more time to accrue citations. Second, the self-selection bias refers to the phenomenon that better papers are selected to be published in preprint repositories in addition to peer-reviewed venues. Moed (2007) states that the self-selection bias can occur due to two reasons: i) the author is very well-reputed and is confident about putting out unfinished work (established author bias), ii) the author, irrespective of their reputation in the field, publish their better works on ArXiv (quality bias). These papers in some cases can be unfinished and the authors keep revising them with time. Controlling for these two biases makes the OA advantage negligible for papers in astrophysics. Since then, there have been multiple studies in different disciplines that study these effects. Davis and Fromerth (2007) find that for mathematics papers there is no significant early view effect and the difference in citations between OA and non-OA can be explained by the self-selection bias. More specifically, they find that there is a quality differential between papers that get submitted to ArXiv compared to papers that do not.
@@
We explore the impact of paper revisions on citations (termed the ”revision effect”). The prior research on this topic is scarce. Rigby, Cox, and Julian (2018) examined the correlation between revisions and the number of citations for a single social sciences journal. We perform a study at a much larger scale for computer science and physics publications and use treatment effect estimation techniques to account for confounders.
@@
The effect of tweeting remains unexplored in computer science and physics. We quantify the impact of tweeting on the number of citations and examine the impact of authors tweeting about their work versus tweets from other accounts.
@@
Differences in publishing trends between computer science and physics (RQ6)
ArXiv’s usage statistics (ArXiv 2024) reveals a significant increase in monthly submissions from January 2014 to March 2024: for physics, from approximately 1,432 to 2,566, and for computer science, from around 1,373 to 8,895, marking a 3.5-fold rise compared to physics. This prompts inquiry into the differing publication dynamics between these fields, including variations in the impact of social promotion. Prior studies have typically relied on RCTs which restricted their scope to individual publication venues or areas, whereas our approach using observational data to calculate effects, allows for the examination of publication dynamics and efects of promotion across computer science and physics.
@@
We create our dataset using data from six sources. In this section, we describe our process of dataset development (detailed in Fig. 2). Then, we describe the distribution of the number of citations of papers belonging to that dataset and how we use the dataset to answer our research questions.
@@
We have created our dataset using data from six sources: Semantic Scholar222www.semanticscholar.org (CC BY-NC license), WikiCFP333www.wikicfp.com, Crossref444www.crossref.org, ArXiv555arxiv.org (CC0: Public Domain license), Altmetric.com, and the X API. Semantic Scholar is a search engine for scientific literature developed by the Allen Institute of AI. We used their publicly available dataset to get information about articles, authors, citations, and highly influential citations. WikiCFP is a website that contains details about calls for papers for scientific and technological conferences. We use the data from WikiCFP as a source of important dates relating to the submission and decision deadlines for conferences in computer science. Similarly, we collect the dates of journal publications for physics papers using the Crossref API. To access information about paper submission and revision dates, we utilize the dataset provided by ArXiv. We examined citations spanning a five-year timeframe from the date of first publication (either on ArXiv or in a conference), incorporating papers up to 2017 in our analysis [^6] Association for Computational Linguistic’s (ACL) anonymity policy started in October 2017 so we have safely ignored that and have not taken the anonymity policy into account in our analysis., as we used Semantic Scholar data up to 2022. Altmetric.com provided us with research access to their proprietary dataset that serves as the data and metadata source for tweets related to each research paper in our dataset. Utilizing the unique tweet ID provided by Altmetric.com, we retrieve the tweet, its metadata, and the author’s metadata using the X API.
@@
Cross-referencing dates for conference and journal publications.
We use WikiCFP as a source of the following 5 dates for computer science conferences: the submission deadline, the notification date (the date when the author is notified about acceptance or rejection), final version (camera-ready) due date, and conference start and end dates. We noticed that a few top conferences and dates for some important conferences in computer science (like NeurIPS, AAAI, and ACL conferences) were not present on WikiCFP. We manually added these missing conferences to our dataset and annotated the missing dates. Details about the annotation process can be found in Appendix B. Despite this, we could not find all 5 required dates for around 150 important conferences, e.g., some of the older NeurIPS editions. Overall, we have 46.5k conference editions with full information about them.
We matched conference names between the WikiCFP and Semantic Scholar datasets to add information about conference dates for the articles on Semantic Scholar by extending the conference name matching algorithm used by Demetrescu et al. (2022), but unlike their approach, we do not rely on DBLP for conference names. Instead, we utilize conference names and acronyms from WikiCFP, and conference names from Semantic Scholar. Appendix C contains the details of our algorithm. Our final dataset after cross-referencing Semantic Scholar, WikiCFP, and ArXiv contains 18,113 articles with author and citation information (including highly influential citations), conference dates, and ArXiv submission data.
For journal articles in physics, we used Crossref as a source for publication dates and cross-referenced this data with papers on Semantic Scholar using the paper’s DOI to create a dataset that contains 501,766 articles.
These datasets are used for calculating trends for ArXiv-first submissions and revisions (RQ1) and to calculate the ArXiv-first and revision effects (RQ2,3).
Figure 3: Comparison of the distribution of citations in 5 years after the first date of publication for all papers on Semantic Scholar (SS), those on ArXiv cross-referenced with Semantic Scholar, and our dataset derived from cross-referencing ArXiv, Semantic Scholar, and WikiCFP/Crossref for (a) computer science and (b) physics. Cit. refers to citations and HI cit. refers to highly influential citations.
@@
The Altmetric.com API provides rich information about papers that are publicized on social media, including fine-grained information about each social media post describing a paper. For computer science papers, using the API, we found 13,564 papers that have at least one tweet associated with them, 2,929 papers that have no tweets associated with them, and Altmetric.com did not contain information about 1,620 papers in our dataset. In physics, 134,563 papers have at least one associated tweet, while 188,692 have no tweets linked to them. Additionally, data for 178,511 papers could not be retrieved using Altmetric.com.
@@
  • Identifying aggregator accounts. To identify aggregators, we set a threshold based on the frequency of tweeting research papers after a close manual inspection, classifying accounts exceeding one paper every 10 days as aggregators. We integrate data on these aggregators’ tweets and followers as a control variable in our analysis.
  • Matching authors names to X profiles. We identify papers that authors tweeted about by cross-referencing author names with X users who tweeted about a paper. We employ a name matching algorithm and estimate false negatives using an independent validation classifier. Our name matching algorithm identifies 1,344 papers tweeted by authors for computer science and 13,812 for physics, with a true positive rate of 94%. We next try to estimate the prevalence of false negatives in our name matching algorithm. Employing a method from Mane, Srivastava, and Hwang (2005), we estimated approximately 170 false negatives for computer science, constituting about 1% of papers in our dataset, which we conclude is negligible. Appendix D contains details on the algorithm and false negative estimation.
@@
Publications in our dataset have a higher number of mean and median citations compared to all computer science or physics papers in Semantic Scholar or ArXiv, since only the higher-tier conferences and journals appear in WikiCFP and Crossref (compare first through third boxes in Fig. 3(a) or  3(b)). Our estimates of treatment effects are for papers that are published in such high-tier conferences and journals. We note that for lower-tier conferences and journals, treatment effects will be proportionally lower. Finally, we notice that the number of influential citations is about 7 times smaller than the number of plain citations for computer science and 20 times smaller for physics (third and fourth box plots in Fig. 3(a) and  3(b)).
We use our entire dataset for computer science and physics to study temporal trends in paper promotion (RQ1). To calculate the effects (RQ2-5), we use all 18,113 papers in computer science and a 10% subsample of the physics dataset which includes 50,177 articles. We opt for this compromise due to resource constraints, as we lacked the capacity to train superlearners containing tree-based models like BART on the entire physics dataset, since these models are not space-efficient and require more than 200 GB of RAM to run a single experiment.
@@
Control Variable Reason to Control
@@
Conference rating (CS)/
Impact Factor of a journal (Physics)